# Boosting Visual Recognition for Autonomous Driving in Real-world Degradations with Deep Channel Prior

### v1

@@ I Introduction
In recent years, significant strides have been witnessed in autonomous driving technology [1][2][3][4][5][6][7][8], primarily attributed to the improved reliability of vehicle environmental perception. In this context, owing to higher performance-cost ratio and installation convenience, the on-board RGB imaging camera has become one of the most widely-used sensor to guarantee the safety of vehicular navigation. However, the image acquisition process in the real world is inevitably corrupted by various factors [9][10][11], including adverse weather conditions, sensor noise, motion blur, and so on, resulting in the reduction in image visibility and the weakness in feature representation capacity of objects. Ultimately, the image quality degradation will lead to declined navigation environment perception [12][13][14][15][16][17], heightening driving risks and raising safety concerns. It is thus necessary to develop available and improved visual recognition methods in real-world degradations.
@@ II Related Work
With the increasing popularity of autonomous vehicles, the safety concerns of autonomous driving are also rising. One of the vital challenges in ensuring the safety of autonomous driving lies in improving the limited performance of its visual recognition system under various degradation scenarios. Fragility of deep neural networks for visual recognition has been studied in recent years [18][27][28][29], which revealed that existing recognition networks trained on HQ images are vulnerable to image distortions such as weather, noise, blur, and low-light. Fine-tuning is a popular strategy to mitigate this issue, however fine-tuning based methods need semantic labels of degraded images to supervise, which are always unavailable in practical scenarios. In addition, data augmentation [30][31] is a simple yet effective method that allows the model to see more distorted images, thereby generalizing well during inference. However, deep models tend to learn an average data distribution when a large amount of data with various degradation types is fed into the networks, resulting in poor performance on each specific type [32]. The other three mainstream methods are as follows:
@@ IV Experiments
This section first discusses the implementation details of our experiments, including datasets, experimental configurations, model fine-tuning, and training settings. To prove the advantages of UFEM for various visual recognition tasks in autonomous driving, we will detailedly compare and analyze the improving of UFEM and other methods for environmental perception across diverse degradation scenarios.
@@ IV-A Dataset and Implementation
Dataset Details. We conduct extensive experiments for evaluating our proposed UFEM on three kinds of high-level vision tasks in autonomous driving, including image classification, object detection, and semantic segmentation. Specifically, we perform evaluations on one large synthetic dataset and seven real-world degradation datasets. Tab. I provides a summary of crucial statistics for the all employed datasets.
Implementation Details. We employ pre-trained VGG16 [53], ResNet50 [54], YOLOv5 [55], and DeepLabv3+ [56] as the baseline models. The UFEM is trained for 200 epochs using the Adam optimizer with a batch size of 5. The initial learning rate of generators and discriminators are set to 2e-4 and 1e-4, respectively. The learning rate is multiplied by 0.1 every 10 epochs. All experiments are run on one Nvidia A100 GPU. More details about the settings of each image restoration (IR) and UDA method can be found in following parts.
TABLE I: Statistics of the datasets used to report results on three different downstream vision tasks, i.e., Image Classification (IC), Object Detection (OD), and Semantic Segmentation (SS). #Volume is the number of images contained in the dataset, #Cls is the number of classes for each dataset.
Dataset Task #Volume #Cls Type Degradation
@@
IV-B Model Fine-tuning and Method Comparison Details
Model Fine-tuning on Clear Dataset. As mentioned above, we conduct comparisons on both synthetic and real datasets. Before conducting comparisons on real datasets, since the categories of Haze-20 [18], ExDARK [27], RTTS [49], and DAWN [57] are inconsistent with the pre-trained models (i.e., VGG16, ResNet50, and YOLOv5), we need to fine-tune the models to obtain the well-performing classifiers and detectors on clear images. For image classification on Haze-20 [18], we first randomly select 100 clear images per class from HazeClear-20 (80 for training, 20 for validation, and the remainder for testing) to fine-tune the pretrained classifiers. After that, we obtain the classifiers for 20 classes, achieving accuracies of 95.9% and 93.9% on the test sets, respectively. For image classification on ExDARK, its fine-tuning is similar to the above, however, there is no clear set corresponding to ExDARK. Therefore, following the collection in [27], we manually select images of the same categories as ExDARK from the ImageNet [61], PASCAL VOC [62], and COCO [63] datasets, filter out the obviously low-light images and ultimately obtain the clear dataset ExDark-clean for fine-tuning. For object detection on RTTS and DAWN, we choose CityScapes [64] as our clear set and filter its categories to be consistent with RTTS (5 classes) and DAWN (6 classes) respectively. After that, fine-tuning is performed on the clear set to obtain a detector that performs well on the clear set. Different from the above, for the ImageNet-C and three nighttime semantic segmentation datasets, whose categories are aligned with the official pre-trained models, we do not fine-tune but choose the pre-trained models as the baselines.
@@ IV-B Model Fine-tuning and Method Comparison Details
TABLE II: Quantitative comparison of our proposed UFEM with Image Restoration (IR) and Unsupervised Domain Adaption (UDA) methods on foggy and motion blur conditions, which contains 5 severity levels.
@@ IV-C Low-Quality Image Classification
Quantitative Results. To fully verify the effectiveness and generality of our UFEM, we compare it with 20 kinds of IR and UDA methods. We conduct the evaluation on one large-scale synthetic dataset (i.e., ImageNet-C [28]) and two real degradation datasets(i.e., Haze-20 [18] and ExDARK [27]). For ImageNet-C, we choose fog and motion blur condition for evaluation. As can be seen from Tab. II and Tab. III, our proposed method significantly surpasses the IR and UDA methods in terms of classification accuracy on synthetic and real degradation datasets. The UDA methods as the general generalization means, struggle to discern specific degradation types, leading to universally lower performance compared to IR methods. In addition, since they rely on unpaired but semantic-aligned images for retraining, which is not suitable for real-world applications, we mainly focus on comparison with the IR methods on real datasets. Continuing, when it comes to real foggy and dark conditions, the improvement of IR methods is also limited, while our method can improve the accuracy of pre-trained networks significantly. For example, on the Haze-20 dataset, the best-performing D4 [66] improves ResNet50 by only 0.3%, while our method can boost by 9.4%. Similar conclusions can be found on the ExDARK dataset. All this demonstrates that our UFEM effectively boosts the performance of pre-trained models, beneficial for autonomous vehicles to improve visual perception and navigational safety under real degradation conditions.
@@
Figure 4: Grad-CAM attention maps comparison of our UFEM with UIR methods. Top-1 prediction of the classifier ResNet50 together with its confidence score is presented with each example. The first two rows are synthetic foggy images from ImageNet-C, and the last two are real dark images from ExDARK.
@@
Grad-CAM Attention Comparison. Besides, we compare the contribution of our UFEM and UIR methods in accurately drawing the attention of classifiers leveraging grad-CAMs [72]. As shown in Fig. 4, most UIR methods fail to effectively guide the classifier to focus more accurately on the discriminative regions of the image. This, in turn, results in misclassifications by the classifier, such as the inability to correctly identify a panda in a tree, a basketball flying in the sky, a dog by the riverside, or a motorcycle being ridden after image-level enhancement. While certain UIR methods (e.g., RefineDNet [20] and D4 [66]) enable the classifiers to correctly identify objects in the enhanced images, their prediction scores are often low, indicating a lack of confidence when dealing with enhanced images as inputs. In contrast, our UFEM better enhances the discriminative regions surrounding the objects, allowing the classifier to attentively focus and predict the correct category accurately. Overall, compared to the image enhancement of UIR methods, the feature correction of our UFEM is more recognition-friendly for deep networks.
T-SNE Visualization. Finally, we leverage t-SNE [50] to visualize the distributions of degraded, clear, and enhanced features to verify the effectiveness of our method in bridging the gap between degraded features and clear ones. Firstly, we present the feature distribution comparisons with other IR methods. Specifically, we select the 8 most common categories from ImageNet as inputs and extract the final layer output of VGG16 as feature embeddings for visualization. As shown in Fig. 6, due to the interference of degradation cues, there is an obvious domain drift between degraded and clear features. Indeed, after enhancement with D4 [66] and SLP [67], some degree of bridging the gap between degraded and clean domains is achieved. However, other IR methods fail to deliver the expected results and may even exacerbate domain shift, as seen with YOLY [42]. In contrast, following enhancement by our UFEM, the distribution between enhanced and clear features is effectively aligned, thus improving the performance of existing models on degraded images.
@@
Figure 6: T-SNE visual comparison of feature embeddings from VGG16 augmented with the proposed UFEM and different UIR methods. Clear images of 8 common categories from ImageNet, and their degraded version Fog-4 are used as model inputs. Finally, the original and the enhanced outputs of VGG16 are fed into t-SNE for visualization. As we can see, our UFEM makes enhanced features have a consistent distribution with clear features, solving the problem of domain drift obviously.
@@ IV-D Foggy Object Detection
Grad-CAM Attention Visualization. We next compare the contribution of our UFEM and the IR methods in accurately drawing the attention of detector leveraging gradCAMs [72]. As shown in Fig. 8, UIR methods (e.g., YOLY [42] and RefineDNet [20]) often fail to effectively guide the detector to focus more accurately on the discriminative regions of the image. This, in turn, leads to missed and false detections by the detector, such as the inability to accurately detect the looming riders in fog and the erroneous focus on background areas like the sky. In contrast, our UFEM better enhances the discriminative regions surrounding the objects, allowing the detector to attentively focus and predict the correct category and spatial location of objects. This demonstrates the recognition-friendly property of our UFEM for deep neural networks.
Object Detection Comparison. We further present visual comparisons of object detection results with YOLOv5 [55]. As shown in Fig. 9, after processing with IR methods, feeding the restored images into the detector occasionally yields positive detection results, as indicated by the green arrows in the images. However, in the majority of cases, it leads to worse detections, as indicated by the red arrows in the images. In certain instances, particularly following D4 [66] enhancement, the detector exhibits a tendency to overlook all objects within the image. In contrast, our UFEM notably mitigates false negatives and false positives in environmental perception, such as cars and motorcycles driving in the haze, and pedestrians on the roadside, thus ensuring the safety and efficacy of vehicle navigation in real degradation scenarios.
Figure 8: Grad-CAM attention maps comparison of our UFEM with UIR methods. It can be seen that after UFEM enhancement, the detector effectively focuses on discriminative object regions within the haze, whereas certain UIR methods, while partially guiding the detector to attend correctly, still exhibit instances of misdirection or omission of attention areas.
Figure 9: Object detection results comparison of our UFEM with IR methods on real foggy object detection datasets. After our UFEM enhancement, the detector can avoid missed and false detections in the haze. However, enhancement with IR methods may lead to worse results. Zoom in for the best view.
TABLE V: Quantitative comparison of our UFEM with IR methods on three nighttime segmentation datasets.
@@
IV-E Nighttime Semantic Segmentation
Quantitative Results. Semantic segmentation relies on the local consistency of the image to perform pixel-level classification for each point [56]. However, the introduction of degradation cues will disrupt the image’s statistical and structural characteristics, posing challenges to pre-trained networks for achieving high-precision semantic segmentation. To confirm the accuracy of UFEM in local feature content restoration, we conduct experiments on three real dark segmentation datasets (i.e., ACDC Nighttime [58], Dark_Zurich [59], and Nighttime Driving [60]), and compare it with six advanced IR methods. As reported in Tab. V, a remarkable observation is that all IR-based methods suffer a significant degradation in performance, rather than improvement. One potential reason is that, existing IR models primarily focus on fixed or lower resolutions for image enhancement [33][38], such as 512×512. They often exhibit poor performance on high-resolution nighttime segmentation images, such as 1920×1080, thereby leading to an accuracy decline. In contrast, our method achieves excellent feature correction across various resolutions, enhancing performance on all three real datasets and exceeding IR methods by 7.5$\sim$10.1%. This further validates the effectiveness and generality of our UFEM.
Figure 10: Segmentation results comparison of our UFEM with IR methods on real nighttime semantic segmentation datasets. It can be seen that after our UFEM enhancement, the segmenter can accurately delineate distant and dense objects in the dark. On the contrary, after being enhanced by the UIR methods, the segmenter may lead to worse segmentation results.
@@ IV-F Ablation Study
In this section, we present comprehensive ablation studies on synthetic and real datasets to demonstrate the effectiveness of the individual components in our method.

### v2

# Boosting Visual Recognition in Real-world Degradations via Unsupervised Feature Enhancement Module with Deep Channel Prior

@@ I Introduction
In recent years, significant strides have been witnessed in autonomous driving technology [1][2][3][4][5][6][7][8][9], primarily attributed to the improved reliability of vehicle environmental perception. In this context, owing to higher performance-cost ratio and installation convenience, the on-board RGB imaging camera has become one of the most widely-used sensor to guarantee the safety of vehicular navigation. However, the image acquisition process in the real world is inevitably corrupted by various factors [10][11][12], including adverse weather conditions, sensor noise, motion blur, and so on, resulting in the reduction in image visibility and the weakness in feature representation capacity of objects. Ultimately, the image quality degradation will lead to declined navigation environment perception [13][14][15][16][17][18], heightening driving risks and raising safety concerns. It is thus necessary to develop available and improved visual recognition methods in real-world degradations.
@@ II Related Work
With the increasing popularity of autonomous vehicles, the safety concerns of autonomous driving are also rising. One of the vital challenges in ensuring the safety of autonomous driving lies in improving the limited performance of its visual recognition system under various degradation scenarios. Fragility of deep neural networks for visual recognition has been studied in recent years [19][28][29][30], which revealed that existing recognition networks trained on HQ images are vulnerable to image distortions such as weather, noise, blur, and low-light. To address this issue, there are currently three mainstream methods as follows:
@@ IV-A Dataset and Implementation
Dataset Details. We conduct extensive experiments for evaluating our UFEM on three kinds of high-level vision tasks in autonomous driving, i.e., image classification (IC), object detection (OD), and semantic segmentation (SS). Specifically, we perform evaluations on one large synthetic dataset and seven real-world degradation datasets. Tab. I provides a summary of crucial statistics for the all employed datasets.
Implementation Details. We employ pre-trained VGG16 [51], ResNet50 [52], YOLOv5 [53], and DeepLabv3+ [54] as the baseline models. The UFEM is trained for 200 epochs using the Adam optimizer with a batch size of 5. The initial learning rate of generators and discriminators are set to 2e-4 and 1e-4, respectively. The learning rate is multiplied by 0.1 every 10 epochs. All experiments are run on one Nvidia A100 GPU. More implementation details on the generator and discriminator structures can be found on our GitHub.
IV-B Model Fine-tuning and Method Comparison Details
Model Fine-tuning on Clear Dataset. As mentioned above, we conduct comparisons on both synthetic and real datasets. Before conducting comparisons on real datasets, since the categories of Haze-20 [19], ExDARK [28], RTTS [47], and DAWN [55] are inconsistent with the pre-trained models, we need to fine-tune the models to obtain the well-performing classifiers and detectors on clear images. For image classification on Haze-20, we first randomly select 100 clear images per class from HazeClear-20 (80 for training, 20 for validation, and the remainder for testing) to fine-tune the pretrained classifiers. After that, we obtain the classifiers for 20 classes, achieving accuracies of 95.9% and 93.9% on the test sets, respectively. For image classification on ExDARK, its fine-tuning is similar to the above, however, there is no clear set corresponding to ExDARK. Therefore, following the collection in [28], we manually select images of the same categories as ExDARK from the ImageNet [56], PASCAL VOC [57], and COCO [58] datasets, filter out the obviously low-light images and ultimately obtain the clear dataset ExDark-clean for fine-tuning. For object detection on RTTS and DAWN, we choose CityScapes [59] as our clear set and filter its categories to be consistent with RTTS (5 classes) and DAWN (6 classes) respectively. After that, fine-tuning is performed on the clear set to obtain a detector that performs well on the clear set. Different from the above, for the ImageNet-C and three nighttime semantic segmentation datasets, whose categories are aligned with the official pre-trained models, we do not fine-tune but choose the pre-trained models as the baselines.
TABLE I: Statistics of the datasets used to report results on three different downstream vision tasks.
Dataset Task Volume Class Type Degradation
@@ IV-B Model Fine-tuning and Method Comparison Details
TABLE II: Quantitative comparison of our proposed UFEM with Image Restoration (IR) and Unsupervised Domain Adaption (UDA) methods on foggy and motion blur conditions, which contains 5 severity levels. The Top-1 and Top-2 performances are marked in bold and bold, respectively. Adding shades of gray are the performances of our UFEM. These settings apply to all subsequent tables.
@@ IV-C Low-Quality Image Classification
Quantitative Results. To fully verify the effectiveness and generality of our UFEM, we compare it with 20 kinds of IR and UDA methods. We conduct the evaluation on one large-scale synthetic dataset (i.e., ImageNet-C [29]) and two real degradation datasets(i.e., Haze-20 [19] and ExDARK [28]). For ImageNet-C, we choose fog and motion blur for evaluation. As can be seen from Tab. II and Tab. III, our proposed method significantly surpasses the IR and UDA methods in terms of classification accuracy on synthetic and real degradation datasets. The UDA methods as the general generalization means, struggle to discern specific degradation types, leading to universally lower performance compared to IR methods. In addition, since they rely on unpaired but semantic-aligned images for retraining, which are not available in real-world applications, we mainly focus on comparison with the IR methods on real datasets. Continuing, when it comes to real foggy and dark conditions, the improvement of IR methods is also limited, while our method can improve the accuracy of pre-trained networks significantly. For example, on the Haze-20 dataset, the best-performing D4 [64] improves ResNet50 by only 0.3%, while our method can boost by 9.4%. Similar conclusions can be found on the ExDARK dataset. All this demonstrates that our UFEM effectively boosts the performance of pre-trained models, beneficial for autonomous vehicles to improve visual perception and navigational safety under real degradation conditions.
@@
Figure 5: Grad-CAM attention maps comparison of our UFEM with UIR methods. We choose the Top-3 IR methods for visualization. Top-1 prediction of the classifier ResNet50 together with its confidence score is presented.
Grad-CAM Attention Comparison. Besides, we compare the contribution of our UFEM and UIR methods in accurately drawing the attention of classifiers leveraging grad-CAMs [70]. As shown in Fig. 5, most UIR methods fail to effectively guide the classifier to focus more accurately on the discriminative regions of the image. This, in turn, results in misclassifications by the classifier, such as the inability to correctly identify a panda in a tree, and a dog by the riverside after image-level enhancement. While certain UIR methods (e.g., RefineDNet [21]) enable the classifiers to correctly identify objects in the enhanced images, their prediction scores are often low, indicating a lack of confidence when dealing with enhanced images as inputs. In contrast, our UFEM better enhances the discriminative regions around objects, enabling focused attention and accurate prediction, thereby benefiting deep network recognition.
T-SNE Visualization. Finally, we leverage t-SNE [48] to visualize the distributions of degraded, clear, and enhanced features to verify the effectiveness of our method in bridging the gap between degraded features and clear ones. Firstly, we present the feature distribution comparisons with other IR methods. Specifically, we select the 8 most common categories from ImageNet and Fog-4 as inputs, and extract the final layer output of VGG16 as feature embeddings for visualization. As shown in Fig. 6, due to the interference of degradation cues, there is an obvious domain drift between degraded and clear features. Indeed, after enhancement with D4 [64] and SLP [65], some degree of bridging the gap between degraded and clean domains is achieved. However, the domain shift still persists. In contrast, following enhancement by our UFEM, the distribution between enhanced and clear features is effectively aligned, improving the performance of existing models on degraded images.
@@
Figure 6: T-SNE visual comparison of feature embeddings from VGG16 with our UFEM and Top-3 UIR methods. Zoom in for the best view.
@@ IV-D Foggy Object Detection
Grad-CAM Attention Visualization. We next compare the contribution of our UFEM and the best-performing IR method in accurately drawing the attention of detector leveraging gradCAMs [70]. As shown in Fig. 8, SLP [65] sometimes fail to effectively guide the detector to focus more accurately on the discriminative regions of the image. This, in turn, leads to missed and false detections by the detector, such as the inability to accurately detect the looming riders in fog and the erroneous focus on background areas like the sky. In contrast, our UFEM better enhances the discriminative regions surrounding the objects, allowing the detector to attentively focus and predict the correct category and spatial location of objects. This demonstrates the recognition-friendly property of our UFEM for deep neural networks.
Object Detection Comparison. We further present visual comparisons of object detection results with YOLOv5 [53]. As shown in Fig. 9, after processing with IR methods, feeding the restored images into the detector occasionally yields positive detection results, as indicated by the green arrows in the images. However, in the majority of cases, it leads to worse detections, as indicated by the red arrows in the images. In certain instances, particularly following D4 [64] enhancement, the detector exhibits a tendency to overlook all objects within the image. In contrast, our UFEM notably mitigates false negatives and false positives in environmental perception, such as cars driving in the haze, thus ensuring the safety of vehicle navigation in real degradation scenarios.
Figure 8: Grad-CAM attention maps comparison of UFEM with UIR methods.
Figure 9: Object detection results comparison of our UFEM with UIR methods.
IV-E Nighttime Semantic Segmentation
Quantitative Results. Semantic segmentation relies on the local consistency of the image to perform pixel-level classification for each point [54]. However, the introduction of degradation cues will disrupt the image’s statistical and structural characteristics, posing challenges to pre-trained networks for achieving high-precision semantic segmentation. To confirm the accuracy of UFEM in local feature content restoration, we conduct experiments on three real dark segmentation datasets (i.e., ACDC Nighttime [60], Dark_Zurich [61], and Nighttime Driving [62]), and compare it with six advanced IR methods. As reported in Tab. V, a remarkable observation is that all IR-based methods suffer a significant degradation in performance, rather than improvement. One potential reason is that existing IR models primarily focus on lower resolutions for restoration [31][36], like 512×512. They may perform poorly on high-resolution nighttime images, like 1920×1080, resulting in reduced accuracy. In contrast, our method achieves excellent feature correction across various resolutions, enhancing performance on all three real datasets and exceeding IR methods by 7.5$\sim$10.1%. This further validates the effectiveness and generality of our UFEM.
TABLE V: Quantitative comparison of our UFEM with IR methods on three nighttime segmentation datasets. Bold indicates the least performance decline, while bold is same as above.
@@ IV-E Nighttime Semantic Segmentation
Figure 10: Segmentation results comparison of our UFEM with UIR methods.